# Single-Image based unsupervised joint segmentation and denoising

### v1

@@ 4.1 Segmentation Step
where $K(u)\coloneqq\nabla u$, $\mathcal{F}(v)\coloneqq\|v\|_{1,2}$ and $\mathcal{G}(u)\coloneqq i_{\mathbb{A}}(u)+\int_{\Omega}(f-\Phi_{\theta_{F}}(f))^{2}u+\int_{\Omega}(f-\Phi_{\theta_{B}}(f))^{2})(1-u)$. It holds that $K:\mathbb{X}\rightarrow\mathbb{Y}$ is a linear mapping between Hilbert spaces $\mathbb{X},\mathbb{Y}$ and $\mathcal{F}:\mathbb{Y}\rightarrow[0,\infty]$ and $\mathcal{G}:\mathbb{X}\rightarrow[0,\infty]$ are convex and lower semi-continuous functionals, i.e. all the necessary assumptions for the application of the primal dual algorithm framework proposed in [10] are fulfilled.
@@ 4.1.1 Discretisation
Its adjoint is given by *$\nabla^{*}(v_{1},v_{2})=\nabla^{*}_{1}v_{1}+\nabla_{2}^{*}v_{2}=:-\operatorname{div}(v_{1},v_{2})$ where $\operatorname{div}\colon\mathbb{H}\times\mathbb{H}\to\mathbb{H}$ is the discrete divergence operator and for $(v_{1},v_{2})\in\mathbb{H}\times\mathbb{H}$ we have
@@
$\lVert\nabla u\rVert_{2,1}\coloneqq\sum_{i}\sqrt{(\nabla_{1}u[i])^{2}+(\nabla_{2}u[i])^{2}}\,.$
The discrete versions of the admissible set and the corresponding indicator function, are $\mathbb{A}=\{u\in\mathbb{H}|0\leq u\leq 1\}$, and $i_{\mathbb{A}}.$ The discretisation of the data fidelity term of energy (5) is written as $\sum_{i}{D}(u[i],\boldsymbol{\theta})$, where
${D}(u,\boldsymbol{\theta})$ $\coloneqq d(u,\theta^{F})+d(1-u,\theta^{B})$ (8)
$d(u,\theta^{F})$ $\coloneqq u\cdot\left(\Phi_{\theta^{F}}(f)-f\right)^{2}$
$d(1-u,\theta^{B})$ $\coloneqq(1-u)\cdot\left(\Phi_{\theta^{B}}(f)-f\right)^{2}.$
@@
$\mathcal{E}_{f,\lambda}^{CV}(u,\boldsymbol{\theta})=i_{\mathbb{A}}(u)+\lambda\lVert\nabla u\rVert_{1,2}+\sum_{i}{D}(u[i],\boldsymbol{\theta})\,.$ (9)
The optimisation problem (9) is in general a non-convex and challenging problem to be solved. We will use alternating minimisation, where we employ for the update step of the segmentation mask $u$ the Chambolle-Pock algorithm [10], while for updating the network parameters $\boldsymbol{\theta}=(\theta^{F},\theta^{B})$, we apply ADAM optimisation [22].
@@ 4.1.2 Segmentation algorithm
We here detail the minimisation of the functional (5) with respect to $u$ for fixed $\boldsymbol{\theta}$, which corresponds to solve problem (7) with
$\mathbb{X}$ $=\mathbb{H}$
$\mathbb{Y}$ $=\mathbb{H}^{2}$
$\mathcal{F}$ $=\lambda\lVert v\rVert_{2,1}$
$K$ =$\nabla$= ∇
$\mathcal{G}$ $=i_{\mathbb{A}}+\sum_{i}{D}(u[i],\boldsymbol{\theta}).$
@@
To implement this algorithm, it is required to compute the Fenchel conjugate $\mathcal{F}^{*}$ of $\mathcal{F}$, as well as the proximal mappings of $\mathcal{F}^{*}$ and $\mathcal{G}$. We start with the derivation of the Fenchel conjugate of $\mathcal{F}$. For $\lVert\cdot\rVert_{2,1}$ it corresponds to the indicator function of the unit ball of the dual norm, resulting in $i_{2,\infty}=\lVert\cdot\rVert^{\ast}_{2,\infty}$. Hence we have *$\mathcal{F}^{*}(v)=i_{2,\infty}(v/\lambda)$, the indicator function of $\{v|\lVert\boldsymbol{v}\rVert_{2,\infty}\leq 1\}\subset(\mathbb{H})^{2}$. As a next step, we compute the proximal operators of $\mathcal{F}^{\ast}$ and $\mathcal{G}$. Recall that the proximal operator of the indicator function $i_{C}$ of some set $C$ is given by the orthogonal projection on $C$. The projection $P_{2,\infty}\colon\mathbb{H}\rightarrow\mathbb{H}^{2}$ onto the unit ball in the $(2,\infty)-$norm is thus obtained by
@@
Further, by introducing $\tilde{f}=\left(f-\Phi_{\theta^{F}}(f)\right)^{2}-\left(f-\Phi_{\theta^{B}}(f)\right)^{2}$, one can finally show that
@@
Algorithm 2 Segmentation algorithm based on the minimisation of the energy functional (5) with respect to $u$ for a fixed $\theta$.
@@ 4.2 Acceleration with a mask prior
$\mathcal{E}^{k}_{f,\lambda}(u,\theta)\coloneqq i_{\mathbb{A}}(u)$ $+\lambda\lvert u\rvert_{\text{TV}}+\int_{\Omega}\left(f-\Phi_{\theta^{F}}(f)\right)^{2}u$ (10)
$+\int_{\Omega}\left(f-\Phi_{\theta^{B}}(f)\right)^{2}(1-u)+\frac{\mu}{2}||u-u_{R}^{k}||^{2}\,.$
@@
Algorithm 3 Alternating optimisation scheme with acceleration.
  Initialise $u^{0}\leftarrow f$ and $\theta^{0}=\theta_{0}$ and $u_{R}^{0}$
@@
      $\boldsymbol{\boldsymbol{\theta}}^{k+1}\leftarrow\operatorname*{argmin}_{\boldsymbol{\theta}}\mathcal{E}^{k}_{f,\lambda}(u^{k+1},\boldsymbol{\theta})$with a few ADAM iterations for $\theta^{F}$ and Chan and Vese update for the background if $\Phi_{\theta^{B}}(f)=\theta^{B}\mathbbm{1}$}
      $u^{k+1}\leftarrow\operatorname*{argmin}_{u}\mathcal{E}^{k}_{f,\lambda}(\boldsymbol{u},\theta^{k})$with Algorithm 4)}
@@
To solve the segmentation problem, we reformulate the optimisation of problem (10) for fixed $\boldsymbol{\theta}$ as $\min_{u}\mathcal{F}(K(u))+\mathcal{G}^{k}(u)$, with
$\mathcal{G}^{k}(u)=i_{\mathbb{A}}(u)+\frac{\mu}{2}||u-u_{R}^{k}||^{2}+\int_{\Omega}\left(f(x)-\Phi_{\theta^{F}}(f)(x)\right)^{2}u(x)+\int_{\Omega}\left(f(x)-\Phi_{\theta^{B}}(f)(x)\right)^{2}(1-u(x)).$
Recalling that $\tilde{f}=\left(f(x)-\Phi_{\theta^{F}}(f)(x)\right)^{2}-\left(f-\Phi_{\theta^{B}}(f)(x)\right)^{2}$, we can show that:
@@
Algorithm 4 Segmentation algorithm based on the minimisation of the energy functional (10) with respect to $u$ for a fixed $\theta$.
@@
  Parameters: $\lambda,\sigma,\tau,\theta$
  Initialisation: $v^{0}\in\mathbb{H}$, $u^{0},\bar{u}^{0}\in\mathbb{H}$
@@ 4.3 Denoising step using Noise2Fast strategy
$\min_{\boldsymbol{\theta}}\sum_{i}{D}(\boldsymbol{u}^{k}[i],\boldsymbol{\theta})=\min_{\boldsymbol{\theta}}\sum_{i}\left(\Phi_{\theta^{F}}(f)[i]-f[i]\right)^{2}\cdot u^{k}[i]+\cdot\left(\Phi_{\theta^{B}}(f)[i]-f[i]\right)^{2}\cdot u_{B}^{k}[i],$ (11)
where $\Phi_{\theta^{F}}$ and $\Phi_{\theta^{B}}$ are (deep) experts respectively dedicated to the denoising of the foreground and background.
We build our denoisers on top of the the Noise2Fast method introduced by Lequyer et al. in [26]. In this paper, the authors propose a fast single image blind denoiser, using a special downsampling strategy. More precisely, their method consists in splitting a given image into smaller parts by using a checkerboard downsampling strategy. From a single image, four images are thus generated, by removing one half of all pixels, and shifting the remaining pixels to fill in the gaps left behind. Then, a network is trained to learn the mappings between the resulting downsampled image pairs. Due to the internal redundancy in form of recurrent patches present in images, and the high degree of self-similarity, the neural network will also be able to denoise the whole image instead of the downsampled ones [4, 46, 18]. For a more detailed description of the Noise2Fast training strategy, such as the network architecture, we refer the reader to [26].
In our approach, we use a different loss function as the one described in the work of Lequyer et al [26]. Instead of considering the whole image domain for training, we restrict the optimisation process for the foreground $\Phi_{\theta^{F}}$ (resp. background $\Phi_{\theta^{B}}$) expert to the current segmentation masks $u^{k}$ (resp. $1-u^{k}$) obtained by Algorithm 2.
@@
the set of training data for $k=0,\dots,N$, with $N$ being the number of iterations of the alternating minimisation.
We then train the two denoising networks, $\Phi_{\theta^{F}}$ and $\Phi_{\theta^{B}}$, restricted to the given regions, $u^{k}$, and $u^{k}_{B}$, i.e. for $(\tilde{f},\tilde{g},\tilde{u},\tilde{u}_{B})\in\mathcal{J}^{k}$ we minimise
$\mathcal{L}_{\boldsymbol{u}}^{k}(\boldsymbol{\theta})=\sum_{i}\left(\Phi_{\theta^{F}}(\tilde{f})[i]-\tilde{g}[i])\right)^{2}\cdot\tilde{u}[i]+\left(\Phi_{\theta^{B}}(\tilde{f})[i]-\tilde{g}([i])\right)^{2}\cdot\tilde{u}_{B}[i].$ (12)
Thus the self-supervised denoisers learn to reconstruct even (resp. odd) lines and columns $\tilde{f}$ of the image thanks to the odd (resp. even) ones $\tilde{g}$. As mentioned above, caused by the self-similarity redundancy, by minimising (12), $\mathcal{L}_{u}^{k}$, we also solve problem (11).
In the next paragraph, we demonstrate the possible applications of three different variants of the proposed joint denoising and segmentation method.
@@ 5 Experiments and Results
The code which was used to obtain the results presented in this work is provided on GitHub (https://github.com/Nadja1611/Single-Image-based-unsupervised-joint-segmentation-and-denoising.git). As a first application, we test our method on the microscopy cell nuclei dataset from the DSB2018 dataset [^1] https://www.kaggle.com/c/data-science-bowl-2018 stemming from the Kaggle 2018 Data Science Bowl challenge. The data consists of a diverse collection of cell nuclei imaged by various fluorescence microscopes. The patches are of size $128\times 128$, and come with manually generated segmentation ground truths. More precisely, we use the noise free data and manually add gaussian noise with three different noise levels, namely 10, 30, and 50. In our experiments, we considered the same subset of images than the one used in [5], where the authors demonstrated that the segmentation of noisy data can be improved by addressing denoising and segmentation in a cooperative (but not fully joint) manner.
@@
We stop our alternating Algorithm 1, as soon as the decrease of energy (5) is less than 15 percent of the previous decrease rate. We tried out a few different strategies, and this one turned out to be the most promising one. We indeed observed that the a criteria based on the change in the energy decay is robust to different scales for the regularisation parameter $\lambda$, and it also adapts to different type of images.
We compare the segmentation performance of our joint denoising and segmentation approach with the convex Chan-Vese model from [12] applied either on the noisy images directly, or on the previously denoised data within a sequential approach. For both the proposed joint approach and the sequential one, we use the same denoised image as starting point for fair comparisons. Further, we test our method against the partially joint denoising and segmentation framework in [5].
@@
We start with the evaluation of our method on a subset of the DSB2018 cell nuclei data which were manually corrupted (noise levels 10, 30 and 50). To this end, we train a foreground denoiser, $\Phi_{\theta^{F}}$, and we assume the background to be constant, i.e. $\Phi_{\theta^{B}}=\theta^{B}\mathbbm{1}.$ For this particular type of images, this assumption is useful, while for images with more structural patterns, this may not be a reasonable choice, and two denoising experts might be necessary.
@@
As a further comparison, we applied the convex Chan-Vese model from [12] directly on the noisy images. The obtained results are depicted in Figures 5 to 7, while the segmentation evaluation metrics are summarised in Table 1. We observe that for all three noise levels, the sequential and Chan-Vese method from [12] struggle with intensity inhomogenities of the cells. These examples highlight the strength of the proposed unified approach, which is capable of segmenting cells with intensities close to the mean value of the background. Notice that the proposed approach does not perform well on the last example due to the presence of intensity inhomogeneities, ascribed to a spatially varying field, the bias field, in the upper left corner of the image. Please note that in this case, evaluating the denoising performance might not be appropriate, as we are assuming a constant background and not applying denoising to the background.
In Table 1, the results obtained by the supervised joint denoising and segmentation method DenoiSeg [5], are summarised. Here, we ran the provided code using 10 training images. More precisely, we used the DSB2018 dataset with noise level zero and added Gaussian noise in the same way as before to all of the 4320 images, among which 3750 were used for training, 670 for validation and the same 50 as we used for our experiments for testing. It has to be mentioned, that for validation all 570 annotated validation images are used in [5], resulting in a total number of 580 annotated images in total during the training process. As displayed in Table 1, this method performs the best. To have a fairer comparison in term of training data, we decided to adapt their method by using 10 images in total (DenoiSeg(10 images in Table 1), 7 for training and 3 for validation. In this setting, all available data are still used for the training of the denoiser, whereas for the segmentation network, the ground truth masks for all but the ten training images are zeroed out. With this smaller level of supervision, our approach outperforms the method of [5].
@@
[…]
@@
noise level $n10$ $n30$ $n50$
baseline 0.820 0.773 0.582
sequential 0.799 0.777 0.735
proposed 0.851 0.825 0.786
DenoiSeg [5] 0.864 0.848 0.818
DenoiSeg (10 images) 0.843 0.820 0.750
Table 1: Dice values obtained on 50 images of the DSB2018 dataset for the compared methods, and three different noise levels. Here, baseline is the convex Chan-Vese [12] method directly applied to the noisy data, while for the sequential method, we first denoise the image using Noise2Fast [26]. Our unsupervised method almost reaches the performance of the fully supervised approach [5].
@@
Figure 8 is a typical example showing the strength of the proposed algorithm compared to intensity-based approaches. In this experiment, we preprocessed the given image of size 256$\times$256 in a way, that both regions have the same mean value, and added gaussian noise as described before, with a noise level of 10. As a consequence, the classical Chan-Vese algorithm totally fails on this example. This model can nevertheless perform well with an adapted hand-crafted transformation of the image to segment. As illustrated in the two last image of Figure 8, when fed with the map of normalized image gradient instead of the original image intensities, the Chan-Vese model is able to segment the two part of the image.
@@
In Figure 9, we display the clean image considered in the experiment of Figure 8, as well as different denoised images with their corresponding quantitative metrics. More precisely, the second image in the figure is obtained by applying the Noise2Fast strategy to the whole image, while the third image is the result of the proposed joint optimisation procedure, where the image is composed using the segmentation mask $u$ and the denoised images from the two denoising experts. Especially in the left region, we can observe a better denoising performance of the proposed method, which is more evident by examining the PSNR (20.36 vs 19.815) and SSIM (0.753 vs 0.696) values.

### v2

@@ 4.1 Segmentation Step
where $K(u)\coloneqq\nabla u$, $\mathcal{F}(v)\coloneqq\|v\|_{1,2},$ and $\mathcal{G}(u)\coloneqq i_{\mathbb{A}}(u)+\int_{\Omega}(f-\mathbf{D}_{F}(f))^{2}u+\int_{\Omega}(f-\mathbf{D}_{B}(f))^{2})(1-u)$. It holds that $K:\mathbb{X}\rightarrow\mathbb{Y}$ is a linear mapping between Hilbert spaces $\mathbb{X},\mathbb{Y}$ and $\mathcal{F}:\mathbb{Y}\rightarrow[0,\infty]$ and $\mathcal{G}:\mathbb{X}\rightarrow[0,\infty]$ are convex and lower semi-continuous functionals, i.e. all the necessary assumptions for the application of the primal dual algorithm framework proposed in [8] are fulfilled.
@@ 4.1.1 Discretisation
Its adjoint is given by $\nabla^{*}(v_{1},v_{2})=\nabla^{*}_{1}v_{1}+\nabla_{2}^{*}v_{2}=:-\operatorname{div}(v_{1},v_{2})$ where $\operatorname{div}\colon\mathbb{H}\times\mathbb{H}\to\mathbb{H}$ is the discrete divergence operator and for $(v_{1},v_{2})\in\mathbb{H}\times\mathbb{H}$ we have
@@
$\lVert\nabla u\rVert_{1,2}\coloneqq\sum_{i}\sqrt{(\nabla_{1}u[i])^{2}+(\nabla_{2}u[i])^{2}}\,.$
The discrete versions of the admissible set and the corresponding indicator function, are $\mathbb{A}=\{u\in\mathbb{H}|0\leq u\leq 1\}$, and $i_{\mathbb{A}}.$ The discretisation of the data fidelity term of energy (5) is written as $\sum_{i}{\boldsymbol{F}}(u[i],\boldsymbol{W})$, where
$\boldsymbol{F}(u,\boldsymbol{W})$ $\coloneqq d(u,F)+d(1-u,B)$ (8)
$d(u,F)$ $\coloneqq u\cdot\left(\mathbf{D}_{F}(f)-f\right)^{2}$
$d(1-u,B)$ $\coloneqq(1-u)\cdot\left(\mathbf{D}_{B}(f)-f\right)^{2}.$
@@
$\mathcal{E}_{f,\lambda}^{CV}(u,\boldsymbol{W})=i_{\mathbb{A}}(u)+\lambda\lVert\nabla u\rVert_{1,2}+\sum_{i}\boldsymbol{F}(u[i],\boldsymbol{W})\,.$ (9)
The optimisation problem (9) is in general a non-convex and challenging problem to be solved. We will use alternating minimisation, where we employ the Chambolle-Pock algorithm [8] for the update step of the segmentation mask $u$, while for updating the network parameters $\boldsymbol{W}=(F,B)$, we apply ADAM optimisation [15].
@@ 4.1.2 Segmentation algorithm
Here, we give details of the minimisation of of the functional (5) with respect to $u$ for fixed $\boldsymbol{W}$. This corresponds to solve problem (7) with
$\mathbb{X}=\mathbb{H},\,\mathbb{Y}=\mathbb{H}^{2},\,\mathcal{F}=\lambda\lVert v\rVert_{1,2},\,K=\nabla,\,\mathcal{G}=i_{\mathbb{A}}+\sum_{i}\boldsymbol{F}(u[i],\boldsymbol{W}).$
@@
To implement this algorithm, it is required to compute the Fenchel conjugate $\mathcal{F}^{*}$ of $\mathcal{F}$, as well as the proximal mappings of $\mathcal{F}^{*}$ and $\mathcal{G}$. We start with the derivation of the Fenchel conjugate of $\mathcal{F}$. For $\lVert\cdot\rVert_{1,2}$ it corresponds to the indicator function of the unit ball of the dual norm, resulting in $i_{2,\infty}=\lVert\cdot\rVert^{\ast}_{2,\infty}$. Hence we have $\mathcal{F}^{*}(v)=i_{2,\infty}(v/\lambda)$, the indicator function of $\{v|\lVert\boldsymbol{v}\rVert_{2,\infty}\leq 1\}\subset\mathbb{H}^{2}$. As a next step, we compute the proximal operators of $\mathcal{F}^{\ast}$ and $\mathcal{G}$. Recall that the proximal operator of the indicator function $i_{C}$ of some set $C$ is given by the orthogonal projection on $C$. The projection $P_{2,\infty}\colon\mathbb{H}\rightarrow\mathbb{H}^{2}$ onto the unit ball in the $(2,\infty)-$norm is thus obtained by
@@
Further, by introducing $\tilde{f}=\left(f-\mathbf{D}_{F}(f)\right)^{2}-\left(f-\mathbf{D}_{B}(f)\right)^{2}$, one can finally show that
@@
Algorithm 2 Segmentation algorithm based on the minimisation of the energy functional (5) with respect to $u$ for a fixed $\boldsymbol{W}$.
@@ 4.2 Acceleration with a mask prior
$\mathcal{E}^{k}_{f,\lambda}(u,\mathbf{D}_{F},\mathbf{D}_{B})\coloneqq$ $i_{\mathbb{A}}(u)+\lambda\lvert u\rvert_{\text{TV}}+\int_{\Omega}\left(f-\mathbf{D}_{F}(f)\right)^{2}u\operatorname{d}\!x$ (10)
$+\int_{\Omega}\left(f-\mathbf{D}_{B}(f)\right)^{2}(1-u)\operatorname{d}\!x+\frac{\mu}{2}||u-u_{R}^{k}||^{2}\,.$
@@
  Input: noisy input image $f$
  Initialisation: $u^{0}\leftarrow f$ and $\boldsymbol{W}^{0}=\boldsymbol{W}_{0}$, $u_{R}^{0}$, choose $\lambda>0$
@@
      $\boldsymbol{\boldsymbol{W}}^{k+1}\leftarrow\operatorname*{argmin}_{\boldsymbol{W}}\mathcal{E}^{k}_{f,\lambda}(u^{k+1},\boldsymbol{W})$with a few ADAM iterations for $F$ and Chan and Vese update for the background if $\mathbf{D}_{B}(f)=\mathbf{D}_{B}\cdot\mathbbm{1}$}
      $u^{k+1}\leftarrow\operatorname*{argmin}_{u}\mathcal{E}^{k}_{f,\lambda}(\boldsymbol{u},\boldsymbol{W}^{k})$with Algorithm 4)}
@@
Algorithm 3 Alternating optimisation scheme with acceleration.
To solve the segmentation problem, we reformulate the optimisation of problem (10) for fixed $\boldsymbol{W}$ as $\min_{u}\mathcal{F}(K(u))+\mathcal{G}^{k}(u)$, with
$\mathcal{G}^{k}(u)=i_{\mathbb{A}}(u)$ $+\frac{\mu}{2}||u-u_{R}^{k}||^{2}+\int_{\Omega}\left(f(x)-\mathbf{D}_{F}(f)(x)\right)^{2}u(x)\operatorname{d}\!x$
$+\int_{\Omega}\left(f(x)-\mathbf{D}_{B}(f)(x)\right)^{2}(1-u(x))\operatorname{d}\!x.$
Recalling that $\tilde{f}=\left(f(x)-\mathbf{D}_{F}(f)(x)\right)^{2}-\left(f-\mathbf{D}_{B}(f)(x)\right)^{2}$, we can show that:
@@
  Initialisation: $v^{0}\in\mathbb{H}$, $u^{0},\bar{u}^{0}\in\mathbb{H},\lambda>0,\sigma>0,\tau>0,\boldsymbol{W}$ fixed
@@
Algorithm 4 Segmentation algorithm based on the minimisation of the energy functional (10) with respect to $u$ for a fixed $\boldsymbol{W}$.
@@ 4.3 Denoising step using Noise2Fast strategy
$\min_{\boldsymbol{W}}\sum_{i}{\boldsymbol{F}}(\boldsymbol{u}^{k}[i],\boldsymbol{W})=\min_{F,B}\sum_{i}\left(\mathbf{D}_{F}(f)[i]-f[i]\right)^{2}\cdot u^{k}[i]+\left(\mathbf{D}_{B}(f)[i]-f[i]\right)^{2}\cdot u_{B}^{k}[i],$ (11)
where $\mathbf{D}_{F}(f)$ and $\mathbf{D}_{B}(f)$ are again the (deep) experts respectively dedicated to the denoising of the foreground and background.
We build our denoisers on top of the Noise2Fast method introduced by Lequyer et al. in [19]. The authors propose a fast single image blind denoiser, using a special downsampling strategy. More precisely, their method consists in splitting a given image into smaller parts by using a checkerboard downsampling strategy. From a single image, four images are thus generated, by removing one half of all pixels, and shifting the remaining pixels to fill in the gaps left behind. Then, a network is trained to learn the mappings between the resulting downsampled image pairs. Due to the internal redundancy in form of recurrent patches present in images, and the high degree of self-similarity, the neural network will also be able to denoise the whole image instead of the downsampled ones [4, 37, 13]. For a more detailed description of the Noise2Fast training strategy, such as the network architecture, we refer the reader to [19].
In our approach, we use a different loss function as the one described in the work of Lequyer et al [19]. Instead of considering the whole image domain for training, we restrict the optimisation process for the foreground $\mathbf{D}_{F}(f)$ (resp. background $\mathbf{D}_{B}(f)$) expert to the current segmentation masks $u^{k}$ (resp. $1-u^{k}$) obtained by Algorithm 2.
@@
the training set for $k=0,\dots,N$, with $N$ being the number of iterations of the alternating minimisation.
We then train the two denoising networks, $\mathbf{D}_{F}(f)$ and $\mathbf{D}_{B}(f)$, restricted to the given regions, $u^{k}$, and $u^{k}_{B}$, i.e. for $(\tilde{f},\tilde{g},\tilde{u},\tilde{u}_{B})\in\mathcal{J}^{k}$ we minimise
$\mathcal{L}_{\boldsymbol{u}}^{k}(\boldsymbol{W})=\sum_{i}\left(\mathbf{D}_{F}(\tilde{f})[i]-\tilde{g}[i])\right)^{2}\cdot\tilde{u}[i]+\left(\mathbf{D}_{B}(\tilde{f})[i]-\tilde{g}([i])\right)^{2}\cdot\tilde{u}_{B}[i].$ (12)
Thus the self-supervised denoisers learn to reconstruct even (resp. odd) pixels and $\tilde{f}$ of the image thanks to the odd (resp. even) ones $\tilde{g}$. As mentioned above, caused by the self-similarity redundancy, by minimising (12), $\mathcal{L}_{u}^{k}$, we also solve problem (11).
In the next paragraph, we demonstrate the possible applications of three different variants of the proposed joint denoising and segmentation method. The code is provided on GitHub (https://github.com/Nadja1611/Self2Seg.git).
@@ 5 Experiments and Results
As a first application, we test our method on the microscopy cell nuclei dataset from the DSB2018 dataset [^1] https://www.kaggle.com/c/data-science-bowl-2018 stemming from the Kaggle 2018 Data Science Bowl challenge. The data consists of a diverse collection of cell nuclei imaged by various fluorescence microscopes. The patches are of size $128\times 128$, and come with manually generated segmentation ground truths. More precisely, we use the noise free data and manually add gaussian noise with three different noise levels, namely 10, 30, and 50. In our experiments, we considered the same subset of images as the one used in [5], where the authors demonstrated that the segmentation of noisy data can be improved by addressing denoising and segmentation in a cooperative (but not fully joint) manner.
@@
We stop our alternating Algorithm 1, as soon as the decrease of energy (5) is less than 15 percent of the previous decrease rate. We tried out a few different strategies, and this one turned out to be the most promising one. We indeed observed that a criteria based on the change in the energy decay is robust to different scales for the regularisation parameter $\lambda$, and it also adapts to different type of images.
We compare the segmentation performance of our joint denoising and segmentation approach with the convex Chan-Vese model from [10] applied either on the noisy images directly, or on the previously denoised data within a sequential approach. For both the proposed joint approach and the sequential one, we use the same denoised image as starting point for fair comparisons. Further, we test our method against the two-stage method [7], and the partially joint deep learning based denoising and segmentation framework of [5].
@@
We start with the evaluation of our method on a subset of the DSB2018 cell nuclei data which were manually corrupted (noise levels 10, 30 and 50). To this end, we train a foreground denoiser, $\mathbf{D}_{F}(f)$, and we assume the background to be constant, i.e. $\mathbf{D}_{B}=B\odot\mathbbm{1}.$ For this particular type of images, this assumption is useful, while for images with more structural patterns, this may not be a reasonable choice, and two denoising experts might be necessary.
@@
As a further comparison, we applied the convex Chan-Vese model from [10] directly on the noisy images. The obtained results are depicted in Figures 5 to 7, while the segmentation evaluation metrics are summarised in Table 1. We observe that for all three noise levels, the sequential and Chan-Vese method from [10] struggle with intensity inhomogenities of the cells. These examples highlight the strength of the proposed unified approach, which is capable of segmenting cells with intensities close to the mean value of the background. Notice that the proposed approach does not perform well on the last example due to the presence of intensity inhomogeneities, ascribed to a spatially varying field, the bias field, in the upper left corner of the image. Please note that in this case, evaluating the denoising performance might not be appropriate, as we are assuming a constant background and we do not apply denoising to the background.
Further, we compared to the two-step approach outlined by Cai et al. in [7]. We employed the provided MATLAB code and fine-tuned the parameter $\lambda$ for three distinct noise levels: $n10$ with $\lambda=4.5$, $n30$ with $\lambda=3.5$, and $n50$ with $\lambda=3.2$. Upon analysis of the outcomes, it became evident that this method encounters difficulties with intensity inhomogeneities and tends to yield excessively smoothed results.
In Table 1, the results obtained by the supervised joint denoising and segmentation method DenoiSeg [5], are summarised. Here, we ran the provided code randomly using 10 training images. More precisely, we used the DSB2018 dataset with noise level zero and added Gaussian noise in the same way as before to all of the 4320 images, among which 3750 were used for training, 670 for validation and the same 50 as we used for our experiments for testing. It has to be mentioned that for validation, all 570 annotated validation images are used in [5], resulting in a total number of 580 annotated images in total during the training process. As displayed in Table 1, this method performs the best. To have a fairer comparison in terms of training data, we decided to adapt their method by using 10 images in total (DenoiSeg 10 in Table 1), 7 for training and 3 for validation. In this setting, all available data are still used for the training of the denoiser, whereas for the segmentation network, the ground truth masks for all but the ten training images are zeroed out. With this smaller level of supervision, our approach outperforms the method of [5].
@@
$\begin{split}\mathcal{E}_{f,\lambda}(u,\boldsymbol{W})=i_{\mathbb{A}}(u)&+\lambda\lvert u\rvert_{\text{TV}}\operatorname{d}\!x+\int_{\Omega}\left[K_{\sigma}\ast\left(f-\mathbf{D}_{F}(f)\right)\right]^{2}u(x)\operatorname{d}\!x\\ &+\int_{\Omega}\left[K_{\sigma}\ast\left(f-\mathbf{D}_{B}(f)\right)\right]^{2}(1-u(x))\operatorname{d}\!x\,\end{split}$
@@
Noise Baseline Sequential Proposed TwoStage [7] DenoiSeg [5] DenoiSeg 10
$n10$ 0.820 0.799 0.851 0.829 0.864 0.843
$n30$ 0.773 0.777 0.825 0.800 0.848 0.820
$n50$ 0.582 0.735 0.786 0.758 0.818 0.750
Table 1: Dice values obtained on 50 images of the DSB2018 dataset for the compared methods, and three different noise levels. Here, baseline is the convex Chan-Vese [10] method directly applied to the noisy data, while for the sequential method, we first denoise the image using Noise2Fast [19].
@@
Figure 8 is a typical example showing the strength of the proposed algorithm compared to intensity-based approaches. In this experiment, we preprocessed the given image of size 256$\times$256 in a way, that both regions have the same mean value, and added gaussian noise as described before, with a noise level of 10. As a consequence, the classical Chan-Vese algorithm totally fails on this example. This model can nevertheless perform well with an adapted hand-crafted transformation of the image to segment. As illustrated in the two last images of Figure 8, when fed with the map of normalized image gradient instead of the original image intensities, the Chan-Vese model is able to segment the two part of the image.
@@
In Figure 9, we display the clean image considered in the experiment of Figure 8, as well as different denoised images with their corresponding quantitative metrics. More precisely, the second image in the figure is obtained by applying the Noise2Fast strategy to the whole image, while the third image is the result of the proposed joint optimisation procedure, where the image is composed using the segmentation mask $u$ and the denoised images from the two denoising experts. Especially in the left regions, we can observe a better denoising performance of the proposed method, which is more evident by examining the PSNR (20.36 vs 19.815) and SSIM (0.753 vs 0.696) values.